in a rock concert 2/7 of spectators purchased Rs 1000 tickets while remaining 15000 purchased Rs 500 tickets How many spectators viewed the concert
Answer by nerdybill(6948)

in a rock concert 2/7 of spectators purchased Rs 1000 tickets while remaining 15000 purchased Rs 500 tickets How many spectators viewed the concert
.
Since 2/7 bought Rs1000 tickets, the remaining purchases 5/7 must have purchased Rs500 tickets.
Let x = total number of spectators
then
(5/7)x = 15000
multiply both sides by (7/5):
(7/5)(5/7)x = (7/5)15000
x = (7/5)15000
x = (7)3000
x = 21000 (answer)